OpenStudy (anonymous):

help please What do you do to check whether a number is rational or irrational? Can you use an example please. How does finding the square root of a number compare to finding the cube root of a number?

OpenStudy (igreen):

Rational numbers can be written in simple fractions. Ex: \(\dfrac{1}{2}\) Irrational numbers can't be written as simple fractions. Ex: \(\pi\)

OpenStudy (anonymous):

ok thank you

OpenStudy (igreen):

Finding the square root is finding the number that multiplies itself two times to get the number in the square root. Ex: \(\sqrt{49} = 7\), because \(7\cdot 7 = 49\). Finding the cube root is finding the number that multiplies itself three times to get the number in the cube root. Ex: \(\sqrt[3]{8} = 2\), because \(2 \cdot 2 \cdot 2 = 8\)

OpenStudy (anonymous):

thank you for helping me

OpenStudy (igreen):

No problem!

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.